Dayshift Nov.18

Richard_Coleman (coleman@fsgi03)
Thu, 18 Nov 1999 15:42:00 CST

Shift Summary for day shift, 11.18.99

On Shift: YBH and RickC

Hours of beam delivered at all intensities: 8:00
Hours of beam delivered at nominal intensity: 4:30
Hours of E799 data (special runs not included): 6:00
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift: 15001-15003

Run 15001- HCC ineff. 4 CsI channels in a row (1167,1169,1277,1279)
horizontal beam profile is bad.
Run 15003 - HCC ineff. 4 CsI channels in a row (1167,1169,1277,1279)


========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Mostly E799 data taking, beam oscillates between 4 - 8.5 E 12 for first
4 hours of shift, started with wide horizontal seed position traced to
wrong current in Switchyard quad Q101. One drt_fresh_start & slow Daq
fresh start also. CAB board C2 swapped to fix inefficient HCC channels
above. Many spills are at half intensity due to MI intermittent problem.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15001 07:00 Wide Horizontal beam profile on NM2SEED2

183 07:35 HCC ineff., Reload ET, kfrend, we are not receiving
SEM for last 34 spills, ktrig hung.

15002 07:48 Decide to stop run , hung, fresh start DAQ

15003 08:40 Fresh start slow Daq (Caen_server.tel was not running.

08:40 Horizontal spot size at target solved by setting
switchyard quad Q101 to correct current (supply was
worked on during shutdown to improve regulation...)

10:12 HCC ineff channels: #1167 35.7 % ineff
1169 50
1277 33
1279 63

85 10:27 Stop trigger, Jason H. swaps Cabboard C2 to fix 4
channels above

15004 11:00 start new run

12:30 beam intensity back up to 8-9 E12

========================================================================

Follow up Notes on Previous Shift

Watch CsI 2893 and 2871 EOP and laser stability.

========================================================================
Items to Follow up on Next Shift

DO NOT UPDATE FERA PEDESTALS. (Valeri)